OpenStudy (anonymous):

loretta beasly purchases a bicycle which regularly sells for $365.00. the bicycle is on sale for $299.00. how much is the markdown?

OpenStudy (anonymous):

18.1%

OpenStudy (anonymous):

reduction is 365 - 299 = 66 markdown (in %age ) = 66/365 * 100% = 18.1%

OpenStudy (anonymous):

if u hv to give the answer in terms of $ then it is $66